In its natural condition a sample has a mass of 1.980 kg and a volume of 0.001 m3 after being completely dried in an oven the mass of the sample is 1.800 kg specific gravity is 2.7 unit weight of water is 10 KN/m3. The degree of sat weuration of soil is
Answer #1

Weight of natural soil =1.98 kg ,

volume of soil =.001 m3 .

Weight of dried soil = 1.8 kg

So , unit weight of soil ,

yt =(1.8/.001) kg /m3

=1.8 *10 kN/m3

=18 kN/m3

Unit weight of water =yw=10 kN/m3

G=specific gravity of soil=2.7

S= degree of saturation =????

e =void ratio

w =water contain= (weight of water /Weight soil )=((1.98-1.8)/1.8)=.1

We know,

YE - Gthe lte - G+ WG telow i Grow) - 1987) Grow) 10 19:5 = 2-7 (1101) jou - e - 4 sain se=WG le again - S - WG OLX.2 7 230-5 so,Degree of saturation is =54%